# Talk2PC: Enhancing 3D Visual Grounding through LiDAR and Radar Point Clouds Fusion for Autonomous Driving

### v1

@@ IV-B Quantitative Results
The comparison between CenterPoint-based prediction head and our proposed C3D-RECHead. As shown in Table VI, we comprehensively evaluate the accuracy of 3D object detection and orientation angle estimation using two different detection head paradigms on the Talk2Radar dataset. Our findings indicate that the proposed C3D-RECHead exhibits a significant performance advantage over CenterPoint, achieving notable improvements in both mAP and mAOS.
[…]
Prompt Motion Depth Velocity
Methods Car Pedes Cyclist Car Pedes Cyclist Car Pedes Cyclist
[…]
BACA 38.9 14.1 21.2 49.7 25.7 42.7 40.8 28.6 32.0
The comparison on the architecture of models. As Fig. 9 shows, we have compared T-RadarNet and TPCNet, with the models taking radar and LiDAR as single-modal inputs, respectively, to analyze the effectiveness of the model architecture. Our results show that, regardless of whether radar or LiDAR is used as the input, TPCNet consistently outperforms T-RadarNet in localization accuracy across all object categories and at various distance intervals from the ego vehicle. This demonstrates the architectural superiority of TPCNet over T-RadarNet.
Statistics of object prediction accuracy at different depths. We partition the predicted objects into six depth intervals, each spanning 10 meters relative to the ego vehicle. As shown in Table VII, the fusion of LiDAR and radar significantly enhances object detection accuracy, particularly for distant objects. The fused model outperforms both LiDAR-only and radar-only modalities, demonstrating the necessity of multi-modal fusion.
@@ IV-C Ablation Experiments
Ablation studies of vital modules in TPCNet. As shown in TABLE VIII: (1) We first conduct an ablation study on the individual components of the DGGF module. Our analysis revealed that dynamic graphs achieve higher accuracy in scene object recognition when compared with static graphs, indicating that dynamic graphs handle environmental redundancy more effectively. Furthermore, in constructing edges between graph nodes, MaxPool outperforms AvgPool in capturing salient inter-node relationships.
(2) We compare the performance of different text encoders on the 3D visual grounding task. Transformer-based encoders significantly outperform Bi-GRU, demonstrating the efficiency of the attention mechanisms in extracting meaningful textual information. However, PointCLIP, which leverages contrastive learning between point clouds and text, outperforms RoBERTa and ALBERT, highlighting its advantage in embedding textual information with 3D point cloud-conditioned features, making it particularly suitable for 3D point cloud-based visual grounding tasks.
@@
Figure 8: Visualization of predicted results by our proposed TPCNet in Talk2Radar dataset.The first row presents the prediction results from the view of camera while the second row shows the results from Bird’s Eye View (BEV).
Comparison of performances on cross attention between LiDAR and radar. TABLE IX shows the results of ablation studies in the BACA module. We compared two model configurations: one where radar features were used as the query while LiDAR features served as the key and value, and the other where LiDAR was used as the query while radar was not incorporated as the key or value. Our findings indicate that when radar served as the primary information source, supplemented by the fine-grained 3D features of LiDAR, TPCNet achieves superior performance in the prompts containing motion and velocity information. However, for samples with depth-related prompts, the fusion method that uses LiDAR as the query outperforms the radar-query configuration. Among the three approaches, BACA demonstrates the best performance across all prompt types. This result highlights the effectiveness of our proposed BACA module in leveraging bidirectional fusion to enrich the environmental context features required by each modality, thereby enhancing its ability to accurately identify the object corresponding to the textual prompt.
Figure 9: Prediction results by LiDAR-only, radar-only and the fusion of LiDAR and radar based on the TPCNet model.
@@ IV-D Visualization and Discussion
Fig. 8 illustrates the prediction results of TPCNet on the Talk2Radar dataset. As shown, TPCNet effectively captures the objects of varying sizes corresponding to textual prompts across different scenarios, even in complex point cloud environments with clutter. The model demonstrates robust performance in detecting both single and multiple objects, as well as objects at varying distances, including those in close proximity and farther away.
Fig. 9 shows the performance of TPCNet under three settings: LiDAR-only, radar-only, and fusion of both sensors. For the first sample, we observe that the radar-based prediction produces false positive detections, while the LiDAR-based model yields accurate predictions but lacks the level of refinement achieved by sensor fusion. In the second-row sample, our proposed fusion method consistently outperforms the two single-sensor approaches. For the final sample, both the LiDAR-only and radar-only models exhibit varying degrees of false positives due to clutter, whereas the fusion-based approach effectively suppresses false positives and accurately localizes the objects.
@@
V Conclusions
This paper focuses on 3D visual grounding for autonomous driving and embodied perception. We pioneeringly propose a novel model, TPCNet, which enables 3D detection of referential objects in a scene by interpreting textual prompts to guide the dual-sensor fusion of LiDAR and radar. In TPCNet, we first introduce the Bidirectional Agent Cross Attention (BACA) module, which efficiently fuses LiDAR and radar features at the minimal cost while preserving the dynamic characteristic paradigm of the attention mechanism, allowing for low-cost, dynamic fusion of sensor features with textual prompt features. Secondly, we propose a dynamic graph-based fusion approach for point clouds and text features, termed Dynamic Gated Graph Fusion (DGGF). DGGF dynamically extracts point cloud regions of interest based on text embeddings and outperforms other fusion methods. Finally, to enhance 3D visual grounding performance and align with the prompts containing object depth information, we propose C3D-RECHead, which achieves superior performance compared to CenterPoint. Overall, TPCNet achieves state-of-the-art 3D visual grounding performance on both the Talk2Radar and Talk2Car datasets, outperforming existing models.

### v2

@@ IV-C Ablation Experiments
Ablation studies of vital modules in TPCNet. As shown in TABLE IX: (1) We first conduct an ablation study on the individual components of the DGGF module. Our analysis reveal that dynamic graphs achieve higher accuracy in scene object recognition when compared with static graphs, indicating that dynamic graphs handle environmental redundancy more effectively. Moreover, in constructing edges between graph nodes, MaxPool outperforms AvgPool in capturing salient inter-node relationships.
(2) We compare the performance of different text encoders on the 3D visual grounding task. Transformer-based encoders significantly outperform Bi-GRU, demonstrating the efficiency of the attention mechanisms in extracting meaningful textual information. However, PointCLIP, which leverages contrastive learning between point clouds and text, outperforms RoBERTa and ALBERT, highlighting its advantage in embedding textual features with point cloud-conditioned features, making it particularly suitable for point cloud-based visual grounding tasks.
@@
Comparison of performances on cross attention between LiDAR and radar. TABLE VIII shows the results of ablation studies in the BACA module. We compared two model configurations: one where radar features were used as the query while LiDAR features served as the key and value, and the other where LiDAR was used as the query while radar was not incorporated as the key or value. Our findings indicate that when radar served as the primary information source, supplemented by the fine-grained 3D features of LiDAR, TPCNet achieves superior performance in the prompts containing motion and velocity information. However, for samples with depth-related prompts, the fusion method that uses LiDAR as the query outperforms the radar-query configuration. Among the three approaches, BACA demonstrates the best performance across all prompt types. This result highlights the effectiveness of our proposed BACA module in leveraging bidirectional fusion to enrich the environmental context features required by each modality, thereby enhancing its ability to accurately identify the object corresponding to the textual prompt.
Figure 10: Prediction by LiDAR-only, radar-only and the fusion of LiDAR and radar based on the TPCNet.
@@ IV-D Visualization and Discussion
Fig. 8 illustrates the prediction of TPCNet on the Talk2Radar dataset. As shown, TPCNet effectively captures the objects of varying sizes corresponding to textual prompts in different scenarios, even in complex environments with clutter. TPCNet demonstrates robust performance in detecting both single and multiple objects, as well as objects at varying distances, including those in close proximity and farther away.
Fig. 10 shows the performance of TPCNet under three settings: LiDAR-only, radar-only, and fusion of both sensors. For the first sample, the radar-based prediction produces false positive detections, while the LiDAR-based model yields accurate predictions but lacks the level of refinement achieved by sensor fusion. In the second-row sample, our proposed fusion method consistently outperforms the two single-sensor approaches. For the final sample, both the LiDAR-only and radar-only models exhibit varying degrees of false positives due to clutter, whereas the fusion-based approach effectively suppresses false positives and accurately localizes the objects.
@@
V Conclusions, Limitations and Future Works
This paper presents TPCNet, a novel model for 3D visual grounding in autonomous driving and embodied perception, leveraging textual prompts to guide dual-sensor fusion of LiDAR and radar. The framework integrates three key modules: (i) the Bidirectional Agent Cross Attention (BACA) module for efficient feature fusion at low computational cost, (ii) the Dynamic Gated Graph Fusion (DGGF) module for adaptive graph-based feature selection, and (iii) the C3D-RECHead module for enhanced localization accuracy using depth-aligned prompts. Collectively, these designs enable TPCNet to achieve state-of-the-art performance on both the Talk2Radar and Talk2Car datasets, surpassing existing approaches.
Despite these advances, several limitations remain. Current 3D visual grounding methods, including TPCNet, cannot fully capture contextual scene information or color attributes due to the absence of RGB image features. This constraint reduces the ability to reason about appearance and fine-grained semantic cues in complex driving environments.
Although dual-sensor systems combining LiDAR and radar can address the majority of challenges in 3D visual grounding, certain queries involving object color still require the integration of a camera, even though this inevitably increases system complexity. In the future, we plan to explore a holographic 3D visual grounding framework composed of three sensors. Such integration is expected to provide more flexible and fine-grained object queries, thereby enhancing robustness and adaptability in autonomous driving scenarios.